What is an equation of a line that is perpendicular to y= 3/5x-2 and passes through point (5,-7)?
Answer by stanbon(75887)

What is an equation of a line that is perpendicular to y= 3/5x-2 and passes through point (5,-7)?
-------
Given line has slope = 3/5
----
Perpendicular line must have slope = -5/3
================================
Form: y = mx + b
Solve for "b":
-7 = (-5/3)*5 + b
-----
-7 = -25/3 + b
b = (-21/3)+(25/3)
b = 4/3
-----
Equation:
y = (-5/3)x + (4/3)
=======================
Cheers,
Stan H.
=================